Solve: dy/dx + ylogy/x = y(logy)^2/x^2
1
Expert's answer
2020-09-22T13:10:37-0400

"\frac{dy}{dx}+\frac{y\log{y}}{x}=\frac{y{\log^2{y}}}{x^2}"

we will seek solution in the form "y=10^{kx}" , "\frac{dy}{dx}={k}\ln{10}*10^{kx}"

"{k}\ln{10}*10^{kx}+10^{kx}\frac{kx}{x}=10^{kx}\frac{(kx)^2}{x^2}"

"{k}{\ln10}+k=k^2"

"k(k-(1+\ln{10})=0"

"k_1=0" ; "k_2=1+{1}\ln{10}"

"y_1=10^{0*x}=1"

"y_2=10^{(1+\ln{10})x}"